OpenStudy (anonymous):

x + y = -4 -3x + 2y = 2

OpenStudy (akashdeepdeb):

Use the substitution method to solve this. x+y = -4 is a given So x = -4 - y Now put this x's value in the second equation and then solve for y. When you get y, put y's value into any of the 2 equations to get x. Understood? :)

OpenStudy (anonymous):

help this graphic don't know how to find the two numbers ( )

OpenStudy (anonymous):

STEP BY STEP 1. make x in the equation x + y =4 subject of formular, that is x= ????? 2 . for the value x in -3x + 2y =2, do not write x, SUBSTITUTE x with the answer you found in the first step. thus -3( ????) + 2y =2 3. solve for y when you get your y, take it to the equation x + y = 4, and instead of y, put the value of y. You have got your x. your answer will be (x,y)